Grade 9Mechanics

Three particles of mass 1kg each are placed at the corners of right angled triangle AOB,O being the origin of the co-ordinate system OA and OB along +ve x-direction and -ve y-direction.position of centre of mass is?

Subramanyam
9 Years agoGrade 9

1 Answer

Vikas TU
9 Years ago
Dear Student,
From the question It is given that
OA =OB = 1m
m1 = m2 = m3 = 1kg
Coordinates of O = (0,0)
Coordinates of A = (1,0)
Coordinates of B = (0,1)
Xcom is given as:
Then  xcm = (m1x1+m2x2+m3x3)/(m1+m2+m3)
                 = 0+1+0/3
                  = 1/3
YCOM is given as:
Then ycm = (m1y1+m2y2+m3y3)/(m1+m2+m3)
                = (0 + 0+1)/3
               = 1/3
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)